# Large Language Models Are Effective Human Annotation Assistants, But Not Good Independent Annotators

### v1

@@ 1 Introduction
We present a case study on ai automation for gtd  [^4] https://www.start.umd.edu/data-tools/GTD (Global Terrorism Database), whose researchers maintain a similar open-source terrorism database used by various organizations. By integrating ai into the event annotation pipeline, we demonstrate that state-of-the-art nlp techniques can reduce human effort in traditional workflows. Specifically, we show that large language models (llms) are far from expert quality annotations, but they can serve as effective assistants for humans to extract event variables, thus reducing the variable coding time of experts.
@@ 3.1 Event Set Curation
An event set is a collection of documents about the same event. Unlike document-level annotations, event details are scattered across multiple documents, each offering partial or complementary information. Annotators must prioritize informative documents (e.g., those with high tf-idf scores), systematically review them to identify unique events, and create event sets that meet predefined criteria and sending them for downstream coding. For example, one document describes a gun attack “in San Carlos City on Sunday, where armed men opened fire on a vehicle in Barangay Palampas,” while another specifies that “the shooting occurred in Barangay Palampas, San Carlos City, Negros Occidental, on February 20,” providing precise details about the incident’s date.
@@ 3.2 Variable Coding
Annotators synthesize information from documents in event set for each event, identifying variables like location and perpetrators [^5] Full list and description in Appendix A.5.. One document describes a violent incident involving “approximately 20 people, some armed with axes, who targeted security guards, workers, and police officers” in “Houston, British Columbia,” while another document states that “far-left anti-pipeline extremists” attacked the “Morice River drill pad site off the Marten Forest Service Road.” Annotators use information from all relevant documents to code variables. These variables include text, enumerated types, and integers, representing diverse data types. We measure coding time with and without llm-extracted variables and calculate the frequency of annotators choosing llm-extracted variables.
@@ A.1 Automated Data Processing
A retrieval model collects news articles from trusted sources. String filters identify documents containing potential terrorism attack information, focusing on keywords such as assault, hostage, and rebel. Then, an algorithm removes duplicates and irrelevant documents from this pool. A Support Vector Machine (svm) uses tf-idf to flag highly relevant documents for manual review.
A.2 Definition of Terrorism and Inclusion Criteria
The cuss team defines a terrorist attack as the threatened or actual use of illegal force and violence by a non-state actor to attain a political, economic, religious, or social goal through fear, coercion, or intimidation. Specifically, an article must have all of the following attributes to be included in the database: {enumerate*}
@@ A.2 Definition of Terrorism and Inclusion Criteria
The perpetrators of the events must be sub-national actors. State-level terrorism is excluded from the database.
@@ A.5 Variable Coding Schema
    Country: the country in which the terrorism event occurred.
@@
    Location: the most specific location (e.g., village name) in which the terrorism event occurred.
@@
    Target: the targeted group of the terrorism event.
@@
    Perpetrator: the group carrying out the terrorism event.
@@
    Kills: Number of people killed during the terrorism event.
@@
    Wounds: Number of people injured during the terrorism event.
@@ A.6 Human-lm Agreement by Variable Type
"The following document describes zero or more terrorism incidents. Segment the document based on incidents mentioned and return an array.

### v2

@@ 1 Introduction
We present a case study on ai automation for gtd  [^4] https://www.start.umd.edu/data-tools/GTD (Global Temporal Events Database), whose researchers maintain a similar open-source temporal event database used by various organizations. By integrating ai into the event annotation pipeline, we demonstrate that state-of-the-art nlp techniques can reduce human effort in traditional workflows. Specifically, we show that large language models (llms) are far from expert quality annotations, but they can serve as effective assistants for humans to extract event variables, thus reducing the variable coding time of experts.
@@ 3.1 Event Set Curation
An event set is a collection of documents about the same event. Unlike document-level annotations, event details are scattered across multiple documents, each offering partial or complementary information. Annotators must prioritize informative documents (e.g., those with high tf-idf scores), systematically review them to identify unique events, and create event sets that meet predefined criteria and sending them for downstream coding. For example, one document describes an incident “in San Carlos City on Sunday, where a men attack a vehicle in Barangay Palampas,” while another specifies that “the incident occurred in Barangay Palampas, San Carlos City, Negros Occidental, on February 20,” providing precise details about the incident’s date.
@@ 3.2 Variable Coding
Annotators synthesize information from documents in event set for each event, identifying variables like location and perpetrators [^5] Full list and description in Appendix A.5.. One document describes an incident involving “approximately 20 people, some with axes, who caused injuries” in “Houston, British Columbia,” while another document states that “far-left anti-pipeline extremists” attacked the “Morice River drill pad site off the Marten Forest Service Road.” Annotators use information from all relevant documents to code variables. These variables include text, enumerated types, and integers, representing diverse data types. We measure coding time with and without llm-extracted variables and calculate the frequency of annotators choosing llm-extracted variables.
@@ A.1 Automated Data Processing
A retrieval model collects news articles from trusted sources. String filters identify documents containing potential attack information, focusing on keywords such as assault, hostage, and rebel. Then, an algorithm removes duplicates and irrelevant documents from this pool. A Support Vector Machine (svm) uses tf-idf to flag highly relevant documents for manual review.
A.2 Definition of An Event and Inclusion Criteria
The cuss team defines an incident as the threatened or actual use of illegal force and violence by a non-state actor to attain a political, economic, religious, or social goal through fear, coercion, or intimidation. Specifically, an article must have all of the following attributes to be included in the database: {enumerate*}
@@ A.2 Definition of An Event and Inclusion Criteria
The perpetrators of the events must be sub-national actors. State-level is excluded from the database.
@@ A.5 Variable Coding Schema
    Country: the country in which the event occurred.
@@
    Location: the most specific location (e.g., village name) in which the event occurred.
@@
    Target: the targeted group of the event.
@@
    Perpetrator: the group carrying out the event.
@@
    Kills: Number of people killed during the event.
@@
    Wounds: Number of people injured during the event.
@@ A.6 Human-lm Agreement by Variable Type
"The following document describes zero or more incidents. Segment the document based on incidents mentioned and return an array.